A certain perfect gas is heated at constant pressure from 15 degrees celcius to 95 degrees celcius, the energy required is 1136 KJ / kg.
When this same gas is heated at constant volume to the same temperatures, the energy required is 808
KJ / kg.
Calculate:
(i) Mass heat at constant pressure, cp
(ii) Mass heat at constant volume, CV
(iii) The value of γ
(iv) The particular gas constant, Rp
(v) Molar mass, Mr
1
Expert's answer
2020-06-11T10:34:50-0400

Solution.

"T_1=15^oC;"

"T_2=95^oC;"

"\Delta Q_1=1136KJ/kg;"

"\Delta Q_2=808KJ/kg;"

"\Delta Q_1=c_p\Delta T\implies c_p=\dfrac{\Delta Q_1}{\Delta T};"

i)"c_p=\dfrac{1136KJ/kg}{80^oC}=14.2KJ/(kg^oC);"

"\Delta Q_2=c_v\Delta T\implies c_v=\dfrac{\Delta Q_2}{\Delta T};"

ii)"c_v=\dfrac{808KJ/kg}{80^oC}=10.1KJ/(kg^oC);"

iii)"\gamma=\dfrac{c_p}{c_v};"

"\gamma=\dfrac{14.2}{10.1}=1.4;"

vi)"R_m=c_p-c_v;"

"R_m=14.2-10.1=4.1KJ\/(kg^oC)" ;

v)"R=R_mM\implies M=\dfrac{R}{R_m};"

"M=\dfrac{8.31}{4100}=0.002kg/mol;"

The gas is hydrogen (H2).

Answer: i)"c_p=14.2KJ\/(kg^oC);"

ii)"c_v=10.1KJ\/(kg^oC);"

iii)"\gamma =1.4;"

iv)"R_m=4.1KJ\/(kg^oC);"

v)"M=0.002kg\/mol."